OpenStudy (anonymous):

f(x) = x3 (8x^2 + 3) f ' (x) = ?

OpenStudy (amistre64):

we can either product rule it or expand and conquer

OpenStudy (amistre64):

e and q is prolly the simplest

OpenStudy (anonymous):

well my homework says to use product rule

OpenStudy (amistre64):

then id go with product rule :)

OpenStudy (anonymous):

but idk how to use the product rule

OpenStudy (anonymous):

8 x^5+3 x^3 = f(x) 40x^4+9x^2=f'(x)

OpenStudy (amistre64):

\[D[fg]=f'g+fg' \]

OpenStudy (amistre64):

r and l work for me, as i tend to want to refer to them as left and right; but they come out a bit backwards \[D[rl]=r'l+rl' \]

OpenStudy (anonymous):

so f ' g stands for basically 8*5 raised to n-1 exponent, and then same for the three?

OpenStudy (amistre64):

close, but lets try this route. f'(x) = [x^3]' (8x^2 + 3) + x^3 (8x^2 + 3)' that might be more intuitive

OpenStudy (amistre64):

x^3' = 3x^2 and 8x^2+3 ' = 16x you agree?

OpenStudy (amistre64):

masc expanded instead of product ruling

OpenStudy (anonymous):

okay, so for the x^3 ' you're substituting x-h - x/ h into it? since its te derivitive

OpenStudy (amistre64):

i could, but since you are taking about product rule I assume you are past the "first principles" and onto the "rules" themselves

OpenStudy (amistre64):

the first principles stuff is to prove the rules; once proved we can just use the rules

OpenStudy (anonymous):

okay, so its X^3 ' * 8x^2?

OpenStudy (anonymous):

product rule = uv' + vu' x3 (8x^2 + 3) u=x3 u'=3x^2 v=8x^2+3 v'= 16x f'(x) = (x3)(16x)+(8x^2+3)(3x^2) simplify

OpenStudy (amistre64):

compter froze :/

OpenStudy (anonymous):

okay gotcha